OpenStudy (anonymous):

Please help me out. :) At what point on the y-axis does the line segment joining (1,1) and (9,7) subtend the greatest angle? (Hint: 1

OpenStudy (loser66):

you have to write out the equation of the line pass through those points

OpenStudy (anonymous):

6x - 8y + 2 = 0 ?

OpenStudy (loser66):

where does it come from?

OpenStudy (anonymous):

i get that on the two points..

OpenStudy (loser66):

I am sorry, I don't know whether it is right or wrong until you show me your work

OpenStudy (anonymous):

i used the formula Y-Y1 = Y2-Y1 / X2-X1 (X-X1)

OpenStudy (loser66):

I think!! (y2-y1)/(x2-x1) =m is slope of the line. Ok, let me try once

OpenStudy (loser66):

\[\frac{y_2-y_1}{x_2-x_1}=\dfrac{7-1}{9-1}=\dfrac{6}{8}=\dfrac{3}{4}\] that is the slope.

OpenStudy (loser66):

now, write out the equation of the line with that slope and passing through the point (1,1) which is y = mx +b \[y = \dfrac{3}{4}x +b\] so, \(b= y-\dfrac{3}{4}x\) now plug (1,1) in \(b = 1-\dfrac{3}{4}=\dfrac{1}{4}\) plug back to the equation y = mx +b \(y = \dfrac{3}{4}x+\dfrac{1}{4}\) May be I am wrong, but I don't know why my result does not look like yours. Please check and tell me what's wrong ?

OpenStudy (loser66):

I believe that the net is so bad so that you "disappear" :) OK!! I will wait until you stay to get help. We will discuss then. :)

OpenStudy (anonymous):